### v2

# A Benchmark for Multi-modal Foundation Models on Low-level Vision: from Single Images to Pairs

Abstract
The rapid development of Multi-modality Large Language Models (MLLMs) has navigated a paradigm shift in computer vision, moving towards versatile foundational models. However, evaluating MLLMs in low-level visual perception and understanding remains a yet-to-explore domain. To this end, we design benchmark settings to emulate human language responses related to low-level vision: the low-level visual perception (A1) via visual question answering related to low-level attributes (e.g. clarity, lighting); and the low-level visual description (A2), on evaluating MLLMs for low-level text descriptions. Furthermore, given that pairwise comparison can better avoid ambiguity of responses and has been adopted by many human experiments, we further extend the low-level perception-related question-answering and description evaluations of MLLMs from single images to image pairs. Specifically, for perception (A1), we carry out the LLVisionQA+ dataset, comprising 2,990 single images and 1,999 image pairs each accompanied by an open-ended question about its low-level features; for description (A2), we propose the LLDescribe+ dataset, evaluating MLLMs for low-level descriptions on 499 single images and 450 pairs. Additionally, we evaluate MLLMs on assessment (A3) ability, i.e. predicting score, by employing a softmax-based approach to enable all MLLMs to generate quantifiable quality ratings, tested against human opinions in 7 image quality assessment (IQA) datasets. With 24 MLLMs under evaluation, we demonstrate that several MLLMs have decent low-level visual competencies on single images, but only GPT-4V exhibits higher accuracy on pairwise comparisons than single image evaluations (like humans). We hope that our benchmark will motivate further research into uncovering and enhancing these nascent capabilities of MLLMs. Datasets will be available at https://github.com/Q-Future/Q-Bench.
Index Terms:
Multi-modality large language models, low-level vision, benchmark, perception, description, assessmentI Introduction
The emergent large language models (LLMs) such as ChatGPT and Bard, as well as their excellent open-source counterparts (e.g., LLaMA [llama], MPT [mpt]), have served as powerful general-purpose assistants, which opens a new era for artificial intelligence (AI) from targeting specific tasks towards general intelligence. Following the advancements of LLMs, multi-modality large language models (MLLMs), as represented by LLaVA [llava], MiniGPT-4 [minigpt4], InstructBLIP [iblip], and Otter [otter], have brought exciting progresses on the vision field as well. They are capable of providing robust general-level abilities on visual perception/understanding and can even seamlessly dialog and interact with humans through natural language. While such abilities of MLLMs have been validated on several vision-language tasks such as image captioning [cococaps], visual question answering [cocovqa], cross-modality grounding [kosmos2], and traditional vision tasks such as image classification or segmentation [lai2023lisa], most attention is paid to the high-level perception and understanding of visual contents. Meanwhile, the ability of MLLMs remains unclear on low-level visual perception and understanding, which play significant roles in image quality assessment (IQA) [koniq, spaq] and its associated tasks on perceiving visual distortions (noises, blurs) [koniqplusplus, wu2023explainable], and other low-level attributes (color, lighting, composition, style, etc) [aadb] that may relate to aesthetics of natural photos [avaiaa] as well as human preferences on emerging computer-graphics generated [zhang2023subjective] or AI-generated images [agiqa3k, imagereward]. These low-level visual abilities are strongly associated with a wide range of applications, such as recommendation [wu2023dover], guidance on camera systems [irpotential], or visual quality enhancement [lpips]. Henceforth, it is crucial to evaluate these general-purpose foundation models in low-level visual perception and understanding, to relieve extensive human resources on giving feedback to every specific low-level task.
In this paper, we propose the first systematic benchmark Q-Bench+ to measure the low-level visual abilities of MLLMs, which is constructed around a key question:
How do MLLMs emulate human ability related to low-level visual perception and understanding?
A basic answer comes from the fundamental capability of MLLMs: vision-conditioned language generation. Specifically, for low-level vision, MLLMs should ideally be able to correctly answer low-level visual questions and precisely describe the low-level information of single images. Henceforth, we define the following two emerging abilities of MLLMs that directly arise from their language generation capability:
Ability 1: Perception of Low-level Attributes. As shown in Fig. 2 (A1-1), like a human, an MLLM should be able to respond accurately to simple questions related to low-level attributes, e.g answering ‘No’ for a blurry image when queried with ‘Is this image clear?’.
Ability 2: Description via Natural Language. As shown in Fig. 2 (A2-1), like a human, an MLLM should be able to describe the quality and other low-level related attributes for single images with natural language. The descriptions should be both complete and accurate.
Although the above two capabilities essentially emulate human perception of low-level vision, they still miss some key capabilities of humans. For example, regarding Fig. 1 (a), some people may consider its clarity to be average, while others may deem it poor, while neither opinion should be considered incorrect; instead, everyone would agree that Fig. 1 (b) is clearer than Fig. 1 (a). On the other hand, for those who regard both Fig. 1 (b) and Fig. 1 (a) as blurry, comparing the clarity between the pair can also provide additional valuable information. Noticing these issues, lots of recent subjective studies [lpips, pieapp, pipal] have adopted the a juxtaposition-based paradigm, that is, collecting human opinions by comparing a pair of images. Based on these insights and recent progresses on MLLMs [emu2, bakllava, geminipro, openai2023gpt4] that officially support more than one images as inputs, we further explore whether MLLMs can similarly emulate respective human capabilities:
Can MLLMs adeptly extract and compare low-level visual information between a pair of images?
On answering this question, we further extend the Perception and Description tasks from single images to image pairs:
Extended Ability 1: Perception of Low-level Attributes for image pairs. As shown in Fig. 2 (A1-2), like a human, an MLLM should be able to respond correctly to low-level questions for image pairs, e.g answering ‘No’ for image pair (first blurrier) when queried with ‘Is the first image clearer?’.
Extended Ability 2: Description via Natural Language for image pairs. As shown in Fig. 2 (A2-2), like a human, an MLLM should be able to describe the similarities (joint information) and differences (comparison) of low-level appearances between a pair of images with natural language.
Despite the direct and concrete abilities above, we also evaluate how MLLMs can perform on the traditional IQA task, a highly abstract task that requires understanding on how the low-level attributes affect human judgements, as follows:
Ability 3: Precise Assessment Aligned with Human Opinions. As depicted in Fig. 2 (A3), an MLLM should be able to predict quantifiable quality scores for images, which can be aligned with the human-rated mean opinion scores (MOS).
To evaluate the three abovementioned abilities, we formulate their respective benchmark settings, as follows:
I-1 LLVisionQA+ Benchmark Dataset
To evaluate the low-level perception ability (A1) on various low-level attributes under diverse circumstances, we construct the LLVisionQA+ dataset, including 2,990 single images and 1,999 image pairs from 10 diverse sources. Aligned with existing practices [mmbench, emabench], each single image or image pair in LLVisionQA+ is equipped with a question, alongside a correct answer and false candidate answers. Specifically, we design three diverse types of questions: Yes-or-No questions, What questions, and How questions. Moreover, we divide low-level concerns for single images into four quadrants, via two axes: (1) distortions (blur, noises, etc) vs other low-level attributes (color, lighting, composition, etc) [atqa]. (2) global perception (e.g., sharpness of the whole picture) vs local content-related in-context perception (e.g., whether the red flower is in focus) [sfa]. On the other hand, we separate the low-level concerns for image pairs into four sub-categories: (1) distortions vs other low-level attributes (similar as above). (2) comparison (e.g., which image is clearer) vs joint analysis (e.g., are both images underexposure). With three types of questions and divided concerns, the proposed LLVisionQA+ dataset provides a holistic benchmark for the low-level perception abilities of MLLMs on both single images and pairs.
I-2 LLDescribe+ Benchmark Dataset
For the description ability (A2), given that the output description is expected to be open-ended, we propose the LLDescribe+ dataset by inviting experts with rich experience in the low-level vision field to write long golden low-level descriptions (average 58 words per description) for 499 single images and 450 image pairs. The long golden low-level descriptions then serve as the reference texts for the single-modal GPT to evaluate MLLM output descriptions. To ensure the evaluation is comprehensive, the quality of MLLM descriptions is evaluated through three dimensions: completeness (punish missing information), preciseness (punish outputs controversial with reference), as well as relevance (punish outputs irrelevant to low-level attributes). With golden descriptions and the multi-dimensional evaluation process participated by GPT, we comprehensively evaluate the low-level description ability of MLLMs.
I-3 IQA Benchmark
For the assessment ability, we utilize plenty of existing IQA databases [koniq, kadid, agiqa3k, zhang2023subjective, spaq, livechallenge] that focus on various low-level appearances of images, to benchmark MLLMs within conventional IQA settings. Specifically, we notice that MLLMs encounter difficulties in providing sufficiently quantifiable outputs, whether instructed to directly rate with texts or provide numerical outputs. To solve this challenge, we propose to extract the softmax pooling result on the logits of the two most frequent tokens (good and poor) under the response template of MLLMs (Fig. 2 (A3)) as their quality predictions. Our studies prove that the proposed softmax-based strategy is generally better correlated with human perception than direct token outputs of MLLMs (via argmax), which bridges between these emergent MLLMs and the traditional IQA task settings. Under this strategy, we evaluate all MLLMs on their precise assessment ability by measuring the correlations between their predictions and human opinion scores in various IQA databases. Furthermore, we propose a prompt-ensemble approach to help boost the IQA performance of MLLMs with the softmax-based strategy.
This work is a substantially extended version of our earlier conference publication [wu2024qbench]. Compared with the conference version, we bring three major changes: (1) Most importantly, we extend the perception and description tasks from single images to image pairs, which provides a more comprehensive benchmark for MLLMs on emulating human low-level visual understanding ability. (2) We update the benchmark with the latest popular MLLMs (evaluated MLLMs increased from 15 to 24), providing a review of the development for MLLMs on low-level vision. (3) We further propose a simple yet effective prompt-ensemble approach, which can help boost the zero-shot performance of MLLMs on the assessment task.
In summary, we systematically explore the potential of MLLMs on three low-level visual abilities: perception, description, and assessment. The three realms compose into the proposed Q-Bench+, a MLLM benchmark on low-level visual tasks. Our contributions can be summarized as three-fold:
We build a benchmark for MLLMs on low-level perception ability. To achieve this, we construct a first-of-its-kind balanced and comprehensive LLVisionQA+ dataset with 2,990 single images and 1,999 image pairs with one low-level-related question-answer pair for each image. The LLVisionQA+ dataset includes three question types and multiple low-level concerns to ensure diversity.
We define a benchmark process to evaluate the low-level description ability of MLLMs, including an LLDescription+ dataset of 499 single images and 450 image pairs with expert-labeled long golden quality descriptions, and a GPT-assisted evaluation to rate MLLM-descriptions in terms of completeness, preciseness, and relevance compared with golden descriptions.
To evaluate precise quality assessment ability, we propose a unified softmax-based quality prediction strategy for all MLLMs based on their probability outputs. Furthermore, we propose a prompt-ensemble approach to help boost the IQA performance of MLLMs with the softmax-based strategy. With its effectiveness validated in our experiments, the proposed strategy sets up a bridge between general-purpose MLLMs and traditional IQA tasks that requires quantifiable scores as outputs.
| Type | Source Dataset | LLVisionQA+ | LLDescribe+ | Full Dataset Size |
| Sampled Size | Sampled Size | for A3 Task | ||
| In-the-wild | KONiQ-10K [koniq] | 600 | 200 | 10,073 |
| SPAQ [spaq] | 800 | 200 | 11,125 | |
| LIVE-FB [paq2piq] | 300 | 50 | 39,810 | |
| LIVE-itw [clive] | 300 | 50 | 1,169 | |
[…]
| AGIQA-3K [agiqa3k] | 198 | 80 | 2,982 | |
| ImageRewardDB [imagereward] | 194 | 29 | excluded in (A3) | |
[…]
| LIVEMultiDistortion [livemultipledistortions] | 15 | 10 | excluded in (A3) | |
| Corrupted COCO [cococaps] | 302 | 50 | excluded in (A3) | |
| Corresponding Task in Q-Bench+ | (A1) Perception | (A2) Description | (A3) Assessment | |
| Benchmark Size (single+pairwise) | 2,990+1,999 | 499+450 | 81,284 | |
II Constructing the Q-Bench+
II-A General Principles
II-A1 Focusing on Low-level Visual Abilities of MLLMs
Unlike existing MLLM benchmarks [seedbench, mmbench, emabench] that aim at all-round abilities, the tasks in Q-Bench+ are constrained with two basic principles: a) Requiring perception and/or understanding on low-level attributes of images; b) Not requiring reasoning (i.e. why) or outside knowledge [okvqa]. We adhere to the principles in designing the perception, description, and assessment tasks, making the proposed Q-Bench+ a focused reflection on the low-level visual abilities of MLLMs.
II-A2 Covering Diverse Low-level Appearances
To cover diverse low-level appearances, we collect multi-sourced images for each task, as depicted in Table I. Among all images in the perception and description tasks, two-thirds are in-the-wild images directly collected from social media posts, smartphones, or professional photography. The rest one-third images are collected after various artificial distortions, or via generative processes (CGI, AIGC). Furthermore, we employ k-means clustering for the low-level attribute indicators to certify that the sub-sampled images retain high diversity. In the assessment task, full images of 7 IQA datasets within all three source types are evaluated through traditional IQA metrics. The diverse and multiple sources of images morph the Q-Bench+ into a holistic and balanced benchmark to fairly evaluate low-level-related abilities.
II-A3 Extending from Single Images to Image Pairs
Evaluating image pairs allows for direct comparison and joint analysis of low-level attributes, which can highlight subtle differences or similarities that might not be evident when images are viewed in isolation. Humans are good at comparing, therefore we believe it is also important to benchmark the low-level visual perception and understanding ability of MLLMs on image pairs. Thus we extend the benchmark (only including single images) in our conference version [wu2024qbench] with image pairs to simulate more complex visual tasks that mirror real-world scenarios and challenge the MLLMs to process and compare multiple visual inputs simultaneously.
II-B Benchmark on Low-level Perception Ability
In the first task of Q-Bench+, we evaluate the low-level perception ability of MLLMs to examine whether they can answer simple natural queries related to low-level attributes. For this purpose, we first collect 2,990 single images (I) from multiple sources (see Table I) with diverse low-level concerns, from which we collect 1,999 image pairs (I’) as well. All image pairs are different from each other but may have one repeated image across different pairs. Then, we collect one low-level-related question (Q), one correct answer to the question (C), and 1-3 candidate false answers (F) for each single image or image pair. The 2,990 (I,Q,C,F) and 1,999 (I’,Q,C,F) tuples compose into the LLVisionQA+ dataset (as illustrated in Fig. 3), the first visual question answering (VQA) dataset in the low-level computer vision field. Specifically, the questions in LLVisionQA+ cover four quadrants of distinct low-level concerns and three question types. After constructing the dataset, the (I,Q,C,F) are together fed into MLLMs for evaluation, while their outputs are further examined by GPT to judge correctness. The details are elaborated as follows.
1) Low-level Visual Concerns for Single Images
Axis 1: Distortions vs Other Low-level Attributes. The primary axis differentiates two categories of low-level perceptual attributes: 1) technical distortions [koniqplusplus], seen as the low-level characteristics that directly degrade the quality of images [paq2piq], and 2) aesthetic-related other low-level attributes [aadb, clipiaa] which are discernible to human perception and evoke varied emotions. Several studies [nima, paq2piq, atqa] follow this paradigm and categorize them through a relative golden standard, that whether the attributes directly improve or degrade picture quality (YesDistortions; NoOthers).
Axis 2: Global Perception vs Local In-context Perception. In recent research on low-level vision, it is observed that human perceptions of low-level visuals often intertwine with higher-level contextual comprehension [sfa, rfugc, wu2022fastervqa, qalign]. For instance, a clear sky might lack complex textures yet display exceptional clarity. Furthermore, localized low-level appearances can deviate from their overall counterparts, as observed by [fastvqa, pvq]. Acknowledging these differences, we curate local in-context perception (Fig. 3 right top) questions, that require MLLMs to grasp the content or other context to answer correctly, while other questions are categorized as global perception (Fig. 3 left top).
2) Low-level Visual Concerns for Image Pairs
Axis 1: Distortions vs Other Low-level Attributes. Same as Axis 1 for single images. Please refer to Sec. 3 1).
Axis 2: Compare vs Joint. This dual approach mimics human visual perception more closely. Humans often use both comparison (looking at differences and similarities) and joint analysis (perceiving images in a unified context) when viewing images. The comparison highlights the differences and similarities between the two images, which is the key component of the full-reference IQA tasks [hore2010image] and other low-level enhancement evaluation tasks [zhang2021no, zhang2022no]. The joint analysis, on the other hand, looks at the images as a combined entity to understand the overall context or to detect patterns that emerge only when the images are considered together.
3) Question Types
In the LLVisionQA+ dataset, we curate three question types, Yes-or-No, What, and How to simulate multiple query forms from humans. The details of the three question types are defined as follows.
Type 1: Yes-or-No Questions. The fundamental type of questions is Yes-or-No, i.e., judgments. Specifically, we notice that some MLLMs especially prefer to respond with yes rather than no. To reduce such biases in our benchmark, though designing questions with answers as yes is easier, we ensure that around 40% of all judgments are with correct answers as no, via querying on contrastive low-level attributes or non-existing low-level attributes.
Type 2: What Questions. Despite Yes-or-No judgments, the what questions are also a common type of queries in recent MLLM benchmarks such as [emabench]. In Q-bench+, they classify low-level attributes in pictures (e.g., What distortion occurs in the image?), or associated context given specific low-level appearances (for in-context perception questions, e.g., Which object in the image is under-exposed?). Unlike Yes-or-No questions, the What questions examine more comprehensive low-level attribute understanding of MLLMs, by requiring correct perception on multiple attributes.
Type 3: How Questions. Despite the two common types, we also include a special type, the How questions, to cover non-extreme appearances [wu2023explainable] of low-level attribute dimensions into our benchmark, as an extension to Yes-or-No questions. As shown in Fig. 3, we can query How is the clarity of the image? for the image with both clear and blurry areas, and answer with Medium. With this special question type, we broaden the Q-bench+ into finer-grained low-level perception.
4) GPT-assisted Evaluation Process
The input query format for MLLMs is as follows:
Single Images:
#User: How is the clarity of the image? (Question)
[IMAGE_TOKEN] (Image)
Choose between one of the following options:
A. High (Correct)B͡. Medium(Wrong)C͡. Low(Wrong)
#Assistant: -
Image Pairs:
#User: Which image is brighter? (Question)
The first image: [IMAGE_TOKEN] (Image 1)
The second image: [IMAGE_TOKEN] (Image 2)
Choose between one of the following options:
A͡. The first image(Wrong) B. The second image(Correct)
#Assistant:
The correct answer has been shuffled and finally uniformly distributed among all choices (A/B/C/D). Moreover, while traditional visual question answering [cocovqa, okvqa] tasks typically employ traditional language metrics (BLEU-4, CIDEr) to compare performance, as observed by recent studies [mplugowl] and validated by us, most MLLMs cannot consistently provide outputs on instructed formats. Given the question above, different MLLMs may reply “A.”, “High”, “The clarity of the image is high.”, “The image is of high clarity.” (all correct), which are difficult to be exhaustively-included under traditional metrics. To solve this problem, we design, validate, and employ a 5-round GPT-assisted evaluation process inspired by [mmbench]. Under this process, the question, correct answers, and MLLM replies are fed into GPT for evaluation.
II-C Benchmark on Low-level Description Ability
In the second task of Q-Bench+, we evaluate the language description ability of MLLMs on low-level information. This task is a sibling task of image captioning [cococaps, flickrcaps, nocaps] that describes image content with natural language, with a specific concern on the low-level appearance of images. To evaluate this ability automatically, we first derive a golden low-level description dataset, denoted as LLDescribe+ (Sec. II-C1), including one long (average 58 words) golden description provided by experts for each of 499 images. With these golden text descriptions, we are able to measure the quality of output low-level descriptions from MLLMs with a single-modal GPT, under the three dimensions: completeness, preciseness, as well as relevance (Sec II-C2). The discussions of the golden descriptions and the evaluation process are as follows.
II-C1 Defining Golden Low-level Descriptions for Images
For the description ability, MLLMs should accurately and completely describe low-level visual information of images. Thus, the ground truths for these MLLMs are also built within a basic principle to cover as many low-level concerns as possible, so long as they are enumerated in Sec. 3 and occur in images. The resulting golden descriptions in LLDescribe+ have an average duration of 58 words, notably longer than common high-level image caption datasets (11 for [nocaps], 10 for [cococaps]). Similar to the LLVisionQA+ dataset for the perception task, the 499 single images and 450 image pairs in LLDescribe+ dataset also include all 10 sources (as in Table I) to cover images with diverse low-level appearances. The golden descriptions on different sources of images are depicted in Fig. 4.
II-C2 Evaluation with Single-modal GPT
After collecting the golden descriptions, we design an input prompt to acquire the output descriptions from MLLMs:
Single Images:
#User: Describe the quality, aesthetics and other low-level appearance of the image in details. (Prompt)
[IMAGE_TOKEN] (Image)
#Assistant: -
Image Pairs:
#User: Compare and jointly analyze the quality, aesthetics and other low-level appearance of the images in details. (Prompt)
The first image: [IMAGE_TOKEN] (Image 1)
The second image: [IMAGE_TOKEN] (Image 2)
#Assistant:
Recent studies [vicuna] have proved single-modal GPT [openai2023gpt4] to be a reliable evaluation tool for pure language tasks. Via the LLDescribe+ dataset, we convert the multi-modality problem into a text-only setting, by matching the MLLM outputs with the golden descriptions with single-modal GPT under three dimensions: (1) Completeness. More matched information with the golden description is encouraged. (2) Preciseness. The controversial information with the golden description is punished. (3) Relevance. More proportions of MLLM outputs should be related to low-level information, instead of others. Each dimension is scored among [0,1,2]. Similar as Sec. 4, we repeat 5 rounds for each single evaluation and collect the weighted average as the final score.
II-D Benchmark on Precise Quality Assessment Ability
| Sub-categories | Question Types | Quadrants of Low-level Concerns | Overall | |||||
[…]
| Distortion | Other | |||||||
| Dev Set / random guess | 50.00% | 27.86% | 33.31% | 37.89% | 38.48% | 38.28% | 35.82% | 37.80% |
[…]
| Emu2-Chat (LLaMA-33B) [emu2] | 71.81% | 67.25% | 56.18% | 64.78% | 63.19% | 63.48% | 72.24% | 65.28% |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 53.33% | 43.70% | 38.00% | 40.81% | 47.40% | 45.45% | 49.23% | 45.05% |
| BakLLava (Mistral-7B) [bakllava] | 66.00% | 56.16% | 51.12% | 51.15% | 61.57% | 53.72% | 72.00% | 57.48% |
| SPHINX [sphinx] | 74.18% | 68.81% | 62.07% | 63.62% | 71.76% | 66.12% | 76.33% | 68.56% |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 72.18% | 57.96% | 56.19% | 56.68% | 69.21% | 53.29% | 72.65% | 61.61% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 66.36% | 58.19% | 50.51% | 49.42% | 65.74% | 54.61% | 70.61% | 58.66% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 65.27% | 64.38% | 56.59% | 56.03% | 67.13% | 61.18% | 67.35% | 62.14% |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 69.45% | 65.27% | 60.85% | 61.67% | 70.14% | 56.91% | 75.10% | 65.35% |
| IDEFICS-Instruct (LLaMA-7B) [idefics] | 56.18% | 44.69% | 44.02% | 42.80% | 54.17% | 44.74% | 56.33% | 48.70% |
| Qwen-VL (QwenLM) [Qwen-VL] | 63.09% | 58.19% | 56.39% | 50.58% | 62.73% | 57.89% | 73.88% | 59.40% |
| Shikra (Vicuna-7B) [shikra] | 65.64% | 47.35% | 49.09% | 48.83% | 59.49% | 50.00% | 64.08% | 54.65% |
| Otter-v1 (MPT-7B) [otter] | 57.09% | 40.71% | 39.55% | 42.22% | 49.31% | 44.08% | 52.65% | 46.35% |
| InstructBLIP (Flan-T5-XL) [iblip] | 67.64% | 59.96% | 55.98% | 56.23% | 65.51% | 58.22% | 69.39% | 61.47% |
| InstructBLIP (Vicuna-7B) [iblip] | 71.64% | 52.65% | 43.81% | 48.64% | 62.50% | 55.59% | 64.90% | 56.72% |
| VisualGLM-6B (GLM-6B) [glm] | 60.18% | 54.20% | 46.25% | 51.75% | 54.40% | 53.62% | 57.14% | 53.78% |
| mPLUG-Owl (LLaMA-7B) [mplugowl] | 66.00% | 54.87% | 44.02% | 51.36% | 55.09% | 54.28% | 65.71% | 55.38% |
| LLaMA-Adapter-V2 [llamaadapterv2] | 66.18% | 59.29% | 52.13% | 57.39% | 56.25% | 63.16% | 64.90% | 59.46% |
| LLaVA-v1 (Vicuna-13B) [llava] | 54.00% | 53.10% | 55.38% | 48.64% | 54.63% | 55.59% | 63.27% | 54.18% |
| MiniGPT-4 (Vicuna-13B) [minigpt4] | 55.82% | 50.22% | 40.37% | 42.02% | 48.38% | 51.97% | 61.22% | 49.03% |
[…]
| Qwen-VL-Max (Close-Source) [Qwen-VL] | 75.60% | 79.43% | 66.09% | 73.39% | 74.08% | 71.0% | 76.92% | 73.63% |
| Gemini-Pro (Close-Source) [geminipro] | 68.80% | 73.74% | 62.34% | 66.30% | 71.34% | 63.91% | 73.09% | 68.16% |
| GPT-4V (Close-Source) [openai2023gpt4] | 76.85% | 79.17% | 67.52% | 73.53% | 76.18% | 72.83% | 76.47% | 74.51% |
| Test Set / random guess | 50.00% | 28.48% | 33.30% | 37.24% | 38.50% | 39.13% | 37.10% | 37.94% |
[…]
| Emu2-Chat (LLaMA-33B) [emu2] | 70.09% | 65.12% | 54.11% | 66.22% | 62.96% | 63.47% | 73.21% | 64.32% |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 62.22% | 35.79% | 36.62% | 41.07% | 49.40% | 45.89% | 49.04% | 45.75% |
| BakLLava (Mistral-7B) [bakllava] | 66.46% | 61.48% | 54.83% | 51.33% | 63.76% | 56.52% | 78.16% | 61.02% |
| SPHINX [sphinx] | 74.45% | 65.50% | 62.13% | 59.11% | 73.26% | 66.09% | 77.56% | 67.69% |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 72.26% | 55.53% | 58.64% | 52.59% | 71.36% | 58.90% | 73.00% | 62.68% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 64.60% | 59.22% | 55.76% | 47.98% | 67.30% | 58.90% | 73.76% | 60.07% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 64.96% | 64.86% | 54.12% | 53.55% | 66.59% | 58.90% | 71.48% | 61.40% |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 68.43% | 62.04% | 61.93% | 56.81% | 70.41% | 57.53% | 77.19% | 64.35% |
| IDEFICS-Instruct (LLaMA-7B) [idefics] | 60.04% | 46.42% | 46.71% | 40.38% | 59.90% | 47.26% | 64.77% | 51.51% |
| Qwen-VL (QwenLM) [Qwen-VL] | 65.33% | 60.74% | 58.44% | 54.13% | 66.35% | 58.22% | 73.00% | 61.67% |
| Shikra (Vicuna-7B) [shikra] | 69.09% | 47.93% | 46.71% | 47.31% | 60.86% | 53.08% | 64.77% | 55.32% |
| Otter-v1 (MPT-7B) [otter] | 57.66% | 39.70% | 42.59% | 42.12% | 48.93% | 47.60% | 54.17% | 47.22% |
| InstructBLIP (Flan-T5-XL) [iblip] | 69.53% | 59.00% | 56.17% | 57.31% | 65.63% | 56.51% | 71.21% | 61.94% |
| InstructBLIP (Vicuna-7B) [iblip] | 70.99% | 51.41% | 43.00% | 45.00% | 63.01% | 57.19% | 64.39% | 55.85% |
| VisualGLM-6B (GLM-6B) [glm] | 61.31% | 53.58% | 44.03% | 48.56% | 54.89% | 55.48% | 57.79% | 53.31% |
| mPLUG-Owl (LLaMA-7B) [mplugowl] | 72.45% | 54.88% | 47.53% | 49.62% | 63.01% | 62.67% | 66.67% | 58.93% |
| LLaMA-Adapter-V2 [llamaadapterv2] | 66.61% | 54.66% | 51.65% | 56.15% | 61.81% | 59.25% | 54.55% | 58.06% |
| LLaVA-v1 (Vicuna-13B) [llava] | 57.12% | 54.88% | 51.85% | 45.58% | 58.00% | 57.19% | 64.77% | 54.72% |
| MiniGPT-4 (Vicuna-13B) [minigpt4] | 60.77% | 50.33% | 43.00% | 45.58% | 52.51% | 53.42% | 60.98% | 51.77% |
[…]
| Qwen-VL-Max (Close-Source) [Qwen-VL] | 73.20% | 81.02% | 68.39% | 70.84% | 74.57% | 73.11% | 80.44% | 73.90% |
| Gemini-Pro (Close-Source) [geminipro] | 71.26% | 71.39% | 65.59% | 67.30% | 73.04% | 65.88% | 73.60% | 69.46% |
| GPT-4V (Close-Source) [openai2023gpt4] | 77.72% | 78.39% | 66.45% | 71.01% | 71.07% | 79.36% | 78.91% | 74.10% |
[…]
| Senior-level Human | 84.31% | 88.94% | 72.02% | 79.65% | 79.47% | 83.90% | 87.07% | 81.74% |
In the third task, we benchmark the ability of MLLMs to provide quantifiable assessment on the overall low-level appearance of images. Unlike the two tasks above, we utilize existing IQA datasets that are collected across a variety of low-level appearances to evaluate how MLLMs can predict quantifiable quality scores aligned with human opinions. All the three types of IQA datasets (in-the-wild, generated, artificially-distorted) as mentioned in Sec. II-A are evaluated, to provide a broad range measurement of the assessment ability of MLLMs. Nevertheless, how to collect quantifiable quality scores from MLLMs remains challenging as their outputs only have weak measurability (Sec. II-D1). Noticing that MLLMs can provide probabilities of tokens, we employ softmax pooling on the logits of good and poor under a simple and direct prompt template, deriving into quantifiable quality scores (Sec. II-D2), as illustrated in Fig. 5. Details as follows.
II-D1 Weak Measurability of MLLM Outputs
In Q-Bench+, we aim to fairly compare the assessment ability between different MLLMs on diverse low-level appearances. Henceforth, our principle is to define a unified, simplest instruction that is applicable for all MLLMs on all IQA datasets. Under this principle, we conduct toy experiments on Shikra [shikra] and LLaVA-v1 [llava], with two simple instruction strategies: (A) Direct Instruction, in which the prompt is designed as simple as “Rate the quality of the image”. The top-frequency answers are good (78%), and poor (20%), with other outputs almost negligible. (B) Numerical Instruction, in which we specifically instruct numerical ratings, with the prompt: “Score the quality of the image from 1 to 5, with 1 as lowest and 5 as highest.”. Under the numerical strategy, the top-frequency answers are 5 (84%), 1 (9%), and 3 (5%); though within the score range, the frequencies of scores 2 and 4 are both less than 1%. The toy experiments imply the weak measurability of MLLM outputs, given that the answers are statistically 1) biased towards positive, 2) biased towards extreme, and 3) with only two effective scales. Therefore, it is necessary to explore extended strategies for MLLMs to provide truly quantifiable outputs for low-level assessment.
II-D2 A Softmax-based Evaluation Strategy
Given the above observations, we design the softmax-based evaluation strategy (Fig. 5) to reduce the negative impacts of the biases and lack of scales. To start with, we design our strategy within the Direct Instruction, which is more general and less biased than the Numerical Instruction. The strategy is based on the observation that two top-frequency outputs, good and poor, can be considered as anchors for better and worse human perception, and the Direct Strategy can be approximated into a binary classification problem on the [SCORE_TOKEN] position, or technically, an argmax between the logits of good () and poor () on this position. In our revised strategy, we modify the argmax into softmax to collect better quantifiable scores:
This simple and generally-applicable strategy enables us to collect quantifiable outputs () from MLLMs with higher correlation to human ratings, as verified in our experimental analysis (Table XII).
II-D3 Prompt Ensemble for Boosting Quantitative Abilities of MLLMs
Multiple synonym prompts can broaden the semantic range, allowing for a more nuanced understanding that might be missed by a single term. Additionally, multiple synonym prompts diminish uncertainty since diverse terms have subtly different meanings, resulting in a more dependable assessment. Specifically, we further choose the combination prompts of [good, high, fine] and [poor, low, bad] to replace good and poor respectively. The quantifiable outputs () can then be altered as:
where indicates the positive token set (from good, fine, high, etc.), while represents the negative token set (from poor, bad, low, etc.). The implementation of the prompt ensemble approach does not add extra computational complexity. The core computation occurs once the input prompt is entered and the language model generates the [SCORE_TOKEN]. After this, we only require tokenization of the words used, followed by the calculation of logits for the [SCORE_TOKEN]. The boosted performance is listed in Table XIII.
| Sub-categories | Question Types | Low-level Concerns | Pairwise Concerns | Overall | ||||
[…]
| Dev Set / random guess | 50.00% | 32.16% | 33.30% | 38.59% | 41.74% | 38.66% | 43.89% | 39.60% |
[…]
| Emu2-Chat (LLaMA-33B) [emu2] | 56.64% | 41.15% | 49.62% | 49.12% | 51.91% | 47.86% | 60.00% | 50.05% |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 68.76% | 33.56% | 38.78% | 46.83% | 54.03% | 47.86% | 55.00% | 49.15% |
| BakLLava (Mistral-7B) [bakllava] | 56.92% | 43.83% | 50.00% | 49.33% | 54.34% | 50.66% | 52.22% | 50.94% |
| mPLUG-Owl2 (Q-Instruct) [q-instruct] | 59.19% | 42.12% | 47.43% | 49.63% | 52.48% | 49.81% | 53.88% | 50.54% |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 58.43% | 39.72% | 48.39% | 49.04% | 51.55% | 47.50% | 60.55% | 49.85% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 60.46% | 42.85% | 41.53% | 47.88% | 51.89% | 46.55% | 59.57% | 49.32% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 56.42% | 42.46% | 48.38% | 48.15% | 53.41% | 48.84% | 54.44% | 49.85% |
[…]
| Qwen-VL-Max (Close-Source) [Qwen-VL] | 71.96% | 62.87% | 65.53% | 69.21% | 62.69% | 67.54% | 66.01% | 67.27% |
| Gemini-Pro (Close-Source) [geminipro] | 64.98% | 51.36% | 54.16% | 58.17% | 56.52% | 57.73% | 57.22% | 57.64% |
| GPT-4V (Close-Source) [openai2023gpt4] | 79.34% | 70.54% | 78.52% | 75.84% | 77.95% | 78.80% | 66.11% | 76.52% |
| Test Set / random guess | 50.00% | 32.03% | 33.16% | 38.95% | 41.95% | 38.69% | 43.70% | 39.82% |
[…]
| Emu2-Chat (LLaMA-33B) [emu2] | 51.94% | 29.78% | 53.84% | 42.01% | 55.71% | 46.26% | 49.09% | 47.08% |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 70.36% | 28.13% | 35.98% | 44.08% | 57.43% | 47.02% | 51.11% | 47.94% |
| BakLLava (Mistral-7B) [bakllava] | 60.09% | 45.42% | 50.86% | 53.09% | 58.82% | 54.52% | 55.55% | 52.75% |
| mPLUG-Owl2 (Q-Instruct) [q-instruct] | 60.24% | 47.46% | 48.78% | 52.81% | 53.97% | 51.42% | 59.11% | 53.15% |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 58.07% | 36.61% | 48.44% | 47.74% | 51.90% | 45.73% | 60.00% | 48.94% |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 60.72% | 42.37% | 50.17% | 49.15% | 59.86% | 52.97% | 49.77% | 52.25% |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 57.34% | 47.45% | 49.13% | 49.01% | 59.51% | 52.06% | 52.00% | 52.05% |
[…]
| Qwen-VL-Max (Close-Source) [Qwen-VL] | 67.65% | 67.56% | 65.35% | 69.09% | 61.18% | 68.65% | 61.29% | 66.99% |
| Gemini-Pro (Close-Source) [geminipro] | 65.78% | 56.61% | 56.74% | 60.42% | 60.55% | 60.46% | 60.44% | 60.46% |
| GPT-4V (Close-Source) [openai2023gpt4] | 79.75% | 69.49% | 84.42% | 77.32% | 79.93% | 81.00% | 68.00% | 78.07% |
[…]
| Senior-level Human | 83.00% | 84.81% | 89.85% | 83.13% | 90.78% | 86.55% | 82.25% | 85.48% |
III Experiment
In Q-Bench+, we evaluate the performance of up to 20 up-to-date popular and competitive open-source as well as 4 close-source commercial MLLMs under zero-shot settings.
III-A Findings on Perception
For a holistic examination of the perception ability of MLLMs, we evaluate the multi-choice correctness of MLLMs on different sub-categories of the LLVision+ dataset, which is equally divided as dev (will be released) and test (will keep private) subsets as shown in Table II and Table III respectively. Only the MLLMs that support multiple images input are included for the perception-pair ability benchmark.
III-A1 Perception for Single Images
a) We are glad that the majority of MLLMs can significantly outperform random guess on all sub-categories as shown in Table II. Considering that all participating MLLMs are without any explicit training on low-level visual attributes, these results show strong potentials for these general-purpose models when further fine-tuned with respective low-level datasets. b) Among all open-source MLLMs, the recently-released SPHINX reaches the best accuracy on this question-answering task, followed by Emu2-Chat and InternLM-XComposer-VL, which show rather close results. By achieving more than 64% accuracy on both subsets, these models show exciting potential as robust low-level visual assistants in the future. c) Another key observation is that almost all methods perceive worse on distortions than other low-level attributes, which indicates that distortion questions are relatively more challenging. d) Close-source MLLMs and Humans. It is widely acknowledged that commercial close-source MLLMs are the leading models in various tasks. To evaluate the low-level perception abilities of these MLLMs, we gauge the accuracy of Qwen-VL-Plus (Alibaba), Qwen-VL-Max (Alibaba), Gemini-Pro (Google), and GPT-4V (OpenAI) on the subsets of LLVision+ dataset. All close-source MLLMs achieve superior performance than all open-source MLLMs on the test subset, which indicates that open-source MLLMs still fall behind on low-level visual ability. GPT-4V exhibits the most competitive performance and outperforms the best open-source MLLM (SPHINX) by a large margin (+6%), and on par accuracy with the Junior-level Human. Despite its prowess, there is still a way to go for GPT-4V before it can match the overall proficiency of the Senior-level Human (with experiences on low-level visual tasks, 7% better than GPT-4V). Furthermore, across all categories, the results show that GPT-4V, much like its open-source counterparts, faces challenges in recognizing distortions.
| Dimensions | Completeness | Precision | Relevance | Sum. | |||||||||
[…]
| InfiMM (Zephyr-7B) [InfiMM] | 29.61% | 62.32% | 7.77% | 0.77 | 29.25% | 31.90% | 38.51% | 1.08 | 2.16% | 22.72% | 74.58% | 1.71 | 3.58 |
| Emu2-Chat (LLaMA-33B) [emu2] | 20.01% | 52.77% | 27.22% | 1.07 | 24.66% | 27.12% | 48.22% | 1.24 | 1.21% | 9.91% | 88.88% | 1.88 | 4.19 |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 25.54% | 61.00% | 13.46% | 0.88 | 41.96% | 32.76% | 25.28% | 0.83 | 2.99% | 11.34% | 85.67% | 1.82 | 3.53 |
| BakLLava (Mistral-7B) [bakllava] | 24.31% | 51.22% | 24.47% | 1.00 | 49.23% | 24.11% | 26.66% | 0.77 | 1.25% | 36.22% | 62.53% | 1.61 | 3.38 |
| SPHINX [sphinx] | 27.96% | 64.36% | 7.33% | 0.79 | 26.16% | 32.42% | 41.01% | 1.14 | 1.69% | 23.00% | 74.61% | 1.72 | 3.65 |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 27.71% | 38.58% | 33.71% | 1.06 | 28.11% | 19.78% | 52.11% | 1.24 | 7.91% | 48.18% | 43.91% | 1.36 | 3.67 |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 27.48% | 54.74% | 17.78% | 0.90 | 30.51% | 26.04% | 43.45% | 1.13 | 10.85% | 60.34% | 28.81% | 1.18 | 3.21 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 27.68% | 53.78% | 18.55% | 0.91 | 25.45% | 21.47% | 53.08% | 1.28 | 6.31% | 58.75% | 34.94% | 1.29 | 3.47 |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 19.94% | 51.82% | 28.24% | 1.08 | 22.59% | 28.99% | 48.42% | 1.26 | 1.05% | 10.62% | 88.32% | 1.87 | 4.21 |
| IDEFICS-Instruct (LLaMA-7B) [idefics] | 28.91% | 59.16% | 11.93% | 0.83 | 34.68% | 27.86% | 37.46% | 1.03 | 3.90% | 59.66% | 36.44% | 1.33 | 3.18 |
| Qwen-VL (QwenLM) [Qwen-VL] | 26.34% | 49.13% | 24.53% | 0.98 | 50.62% | 23.44% | 25.94% | 0.75 | 0.73% | 35.56% | 63.72% | 1.63 | 3.36 |
| Shikra (Vicuna-7B) [shikra] | 21.14% | 68.33% | 10.52% | 0.89 | 30.33% | 28.30% | 41.37% | 1.11 | 1.14% | 64.36% | 34.50% | 1.33 | 3.34 |
| Otter-v1 (MPT-7B) [otter] | 22.38% | 59.36% | 18.25% | 0.96 | 40.68% | 35.99% | 23.33% | 0.83 | 1.95% | 13.20% | 84.85% | 1.83 | 3.61 |
| Kosmos-2 [kosmos2] | 8.76% | 70.91% | 20.33% | 1.12 | 29.45% | 34.75% | 35.81% | 1.06 | 0.16% | 14.77% | 85.06% | 1.85 | 4.03 |
| InstructBLIP (Flan-T5-XL) [iblip] | 23.16% | 66.44% | 10.40% | 0.87 | 34.85% | 26.03% | 39.12% | 1.04 | 14.71% | 59.87% | 25.42% | 1.11 | 3.02 |
| InstructBLIP (Vicuna-7B) [iblip] | 29.73% | 61.47% | 8.80% | 0.79 | 27.84% | 23.52% | 48.65% | 1.21 | 27.40% | 61.29% | 11.31% | 0.84 | 2.84 |
| VisualGLM-6B (GLM-6B) [glm] | 30.75% | 56.64% | 12.61% | 0.82 | 38.64% | 26.18% | 35.18% | 0.97 | 6.14% | 67.15% | 26.71% | 1.21 | 2.99 |
| mPLUG-Owl (LLaMA-7B) [mplugowl] | 28.28% | 37.69% | 34.03% | 1.06 | 26.75% | 18.18% | 55.07% | 1.28 | 3.03% | 33.82% | 63.15% | 1.60 | 3.94 |
| LLaMA-Adapter-V2 [llamaadapterv2] | 30.44% | 53.99% | 15.57% | 0.85 | 29.41% | 25.79% | 44.80% | 1.15 | 1.50% | 52.75% | 45.75% | 1.44 | 3.45 |
| LLaVA-v1 (Vicuna-13B) [llava] | 34.10% | 40.52% | 25.39% | 0.91 | 30.02% | 15.15% | 54.83% | 1.25 | 1.06% | 38.03% | 60.91% | 1.60 | 3.76 |
| MiniGPT-4 (Vicuna-13B) [minigpt4] | 34.01% | 32.15% | 33.85% | 1.00 | 29.20% | 15.27% | 55.53% | 1.26 | 6.88% | 45.65% | 47.48% | 1.41 | 3.67 |
| Dimensions | Completeness | Precision | Relevance | Sum. | |||||||||
[…]
| InfiMM (Zephyr-7B) [InfiMM] | 30.75% | 62.66% | 6.22% | 0.75 | 34.17% | 38.84% | 26.35% | 0.91 | 2.57% | 30.84% | 65.28% | 1.61 | 3.28 |
| Emu2-Chat (LLaMA-33B) [emu2] | 41.25% | 54.33% | 4.42% | 0.63 | 38.11% | 36.41% | 25.48% | 0.87 | 4.12% | 38.61% | 57.27% | 1.53 | 3.03 |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 37.95% | 52.17% | 9.11% | 0.70 | 37.68% | 37.33% | 23.73% | 0.84 | 3.95% | 31.15% | 62.84% | 1.56 | 3.12 |
| BakLLava (Mistral-7B) [bakllava] | 29.46% | 59.77% | 10.57% | 0.80 | 40.0% | 38.08% | 21.33% | 0.80 | 2.26% | 15.06% | 82.04% | 1.79 | 3.40 |
| mPLUG-Owl2 (Q-Instruct) [mplug2] | 15.25% | 65.76% | 18.32% | 1.02 | 39.44% | 40.18% | 19.62% | 0.79 | 0.09% | 9.86% | 89.02% | 1.87 | 3.69 |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 19.43% | 65.54% | 14.45% | 0.94 | 30.94% | 43.71% | 24.63% | 0.92 | 3.79% | 26.94% | 68.28% | 1.63 | 3.50 |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 19.68% | 72.57% | 7.19% | 0.86 | 38.00% | 40.04% | 20.97% | 0.82 | 2.13% | 39.77% | 56.66% | 1.53 | 3.22 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 18.77% | 73.44% | 7.79% | 0.89 | 34.66% | 38.72% | 26.62% | 0.92 | 1.02% | 34.59% | 64.39% | 1.63 | 3.44 |
III-A2 Perception for Image Pairs
Perception for image pairs is far more difficult for MLLMs since this task not only requires MLLMs to have stable low-level visual capabilities, but also requires MLLMs to be able to analyze two images simultaneously and conduct discerning comparisons. To enrich the MLLM diversity, we further include the mPLUG-Owl2 fine-tuned with the single image low-level visual dataset Q-Instruct [q-instruct] for comparison. The performance is exhibited in Table III. With closer inspections, we can obtain several interesting findings. a) Open-source MLLMs are poor low-level comparators. It seems that although they might show strong performance for single image perception, they are quite confused by the image pairs. Most of them get worse performance on the Compare subset than the Joint subset, which further confirms this point. For mPLUG-Owl2 (Q-Instruct), despite being fine-tuned with the single image low-level visual dataset Q-Instruct [q-instruct], the overall performance improvement from the low-level knowledge infusion of single images is relatively weak. This also suggests that there is a necessity to build open-source low-level datasets for multiple images to cultivate the corresponding capabilities of open-source MLLMs. b) Close-source MLLMs are more robust in this task. This may be because these close-source MLLMs are supported by training on multiple-image data, allowing them to make better comparative judgments. Particularly with GPT-4V, its performance in the compare subset is significantly higher than in the joint subset, and it far exceeds all other models, even reaching the level of a junior human. c) Perception for image pairs is easier for humans. Comparing image pairs is simpler for humans, as the answers to related questions tend to be more objective. Especially for junior-level humans with no professional experience, they may have stronger subjectivity in grasping absolute sensations, but it is easier to remain objective when they are faced with comparison concepts. For example, it’s difficult for a junior-level human to judge whether the lighting in a dimly lit single image is appropriate. However, if presented with another image with even weaker lighting, they can easily determine which image is worse. This may explain the notable 6% improvements for junior-level human from single images to image pairs.
In conclusion, the performance of open-source MLLMs on low-level perception for image pairs is still far from satisfactory, which needs to be enhanced and optimized.
III-B Findings on Description
III-B1 Description for single images
For the description ability exhibited in Table IV, InternLM-XComposer-VL reaches the best proficiency, especially in terms of the relevance dimension. Nevertheless, in the perspective of the completeness and precision of the descriptions, even the best of all MLLMs cannot obtain an excellent score; on the contrary, almost all MLLMs reach an acceptable standard (0.8/2.0). In general, all MLLMs at present are only with relatively limited and primary ability to provide low-level visual descriptions.
| Dataset Type | In-the-wild | Generated | Artificial | Average | ||||
[…]
| NIQE [niqe] | 0.316/0.377 | 0.693/0.669 | 0.211/0.288 | 0.480/0.451 | 0.075/0.056 | 0.562/0.517 | 0.374/0.428 | 0.387/0.398 |
| CLIP-ViT-Large-14 [clip] | 0.468/0.505 | 0.385/0.389 | 0.218/0.237 | 0.307/0.308 | 0.285/0.290 | 0.436/0.458 | 0.376/0.388 | 0.354/0.368 |
[…]
| Emu2-Chat (LLaMA-33B) [emu2] | 0.664/0.714 | 0.712/0.698 | 0.355/0.341 | 0.597/0.611 | 0.224/0.269 | 0.759/0.751 | 0.841/0.790 | 0.593/0.596 |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 0.124/0.123 | 0.125/0.179 | 0.164/0.133 | 0.225/0.176 | 0.118/0.116 | 0.368/0.317 | 0.099/0.088 | 0.174/0.161 |
| BakLLava (Mistral-7B) [bakllava] | 0.389/0.390 | 0.406/0.398 | 0.227/0.216 | 0.335/0.337 | 0.179/0.209 | 0.542/0.561 | 0.344/0.361 | 0.346/0.353 |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 0.196/0.252 | 0.589/0.614 | 0.217/0.286 | 0.293/0.342 | -0.024/-0.032 | 0.473/0.492 | 0.541/0.546 | 0.326/0.357 |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 0.463/0.459 | 0.443/0.467 | 0.305/0.321 | 0.344/0.358 | 0.321/0.333 | 0.672/0.738 | 0.417/0.440 | 0.424/0.445 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 0.448/0.460 | 0.563/0.584 | 0.310/0.339 | 0.445/0.481 | 0.285/0.297 | 0.664/0.754 | 0.390/0.400 | 0.444/0.474 |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 0.564/0.615 | 0.730/0.750 | 0.360/0.416 | 0.612/0.676 | 0.243/0.265 | 0.732/0.775 | 0.546/0.572 | 0.541/0.581 |
| IDEFICS-Instruct (LLaMA-7B) [idefics] | 0.375/0.400 | 0.474/0.484 | 0.235/0.240 | 0.409/0.428 | 0.244/0.227 | 0.562/0.622 | 0.370/0.373 | 0.381/0.396 |
| Qwen-VL (QwenLM) [Qwen-VL] | 0.470/0.546 | 0.676/0.669 | 0.298/0.338 | 0.504/0.532 | 0.273/0.284 | 0.617/0.686 | 0.486/0.486 | 0.475/0.506 |
| Shikra (Vicuna-7B) [shikra] | 0.314/0.307 | 0.320/0.337 | 0.237/0.241 | 0.322/0.336 | 0.198/0.201 | 0.640/0.661 | 0.324/0.332 | 0.336/0.345 |
| Otter-v1 (MPT-7B) [otter] | 0.406/0.406 | 0.436/0.441 | 0.143/0.142 | -0.008/0.018 | 0.254/0.264 | 0.475/0.481 | 0.557/0.577 | 0.323/0.333 |
| Kosmos-2 [kosmos2] | 0.255/0.281 | 0.644/0.641 | 0.196/0.195 | 0.358/0.368 | 0.210/0.225 | 0.489/0.491 | 0.359/0.365 | 0.359/0.367 |
| InstructBLIP (Flan-T5-XL) [iblip] | 0.334/0.362 | 0.582/0.599 | 0.248/0.267 | 0.113/0.113 | 0.167/0.188 | 0.378/0.400 | 0.211/0.179 | 0.290/0.301 |
| InstructBLIP (Vicuna-7B) [iblip] | 0.359/0.437 | 0.683/0.689 | 0.200/0.283 | 0.253/0.367 | 0.263/0.304 | 0.629/0.663 | 0.337/0.382 | 0.389/0.446 |
| VisualGLM-6B (GLM-6B) [glm] | 0.247/0.234 | 0.498/0.507 | 0.146/0.154 | 0.110/0.116 | 0.209/0.183 | 0.342/0.349 | 0.127/0.131 | 0.240/0.239 |
| mPLUG-Owl (LLaMA-7B) [mplugowl] | 0.409/0.427 | 0.634/0.644 | 0.241/0.271 | 0.437/0.487 | 0.148/0.180 | 0.687/0.711 | 0.466/0.486 | 0.432/0.458 |
| LLaMA-Adapter-V2 [llamaadapterv2] | 0.354/0.363 | 0.464/0.506 | 0.275/0.329 | 0.298/0.360 | 0.257/0.271 | 0.604/0.666 | 0.412/0.425 | 0.381/0.417 |
| LLaVA-v1 (Vicuna-13B) [llava] | 0.462/0.457 | 0.442/0.462 | 0.264/0.280 | 0.404/0.417 | 0.208/0.237 | 0.626/0.684 | 0.349/0.372 | 0.394/0.416 |
| MiniGPT-4 (Vicuna-13B) [minigpt4] | 0.239/0.257 | 0.238/0.253 | 0.170/0.183 | 0.339/0.340 | 0.252/0.246 | 0.572/0.591 | 0.239/0.233 | 0.293/0.300 |
III-B2 Description for image pairs
We also include the mPLUG-Owl2 (Q-Instruct) for description-pair ability benchmark. As shown in Table V, similarly, all MLLMs perform better in the aspect of relevance than completeness and precision. Furthermore, fine-tuned with the single image low-level visual dataset Q-Instruct [q-instruct], mPLUG-Owl2 (Q-Instruct) achieves the best performance on completeness and relevance but gets the lowest score on precision. This indicates the knowledge infusion from single images can effectively enhance an MLLM to focus on corresponding low-level dimensions for targeted responses, but it does not improve the accuracy of the content, meaning it cannot enhance the core analytical ability of image pairs.
III-C Findings on Assessment
| Dataset Type | In-the-wild | Generated | Artificial | |||||
[…]
| Shikra (Vicuna-7B) [shikra] | argmax | 0.178/0.201 | 0.277/0.281 | 0.152/0.169 | 0.248/0.267 | 0.071/0.065 | 0.513/0.562 | 0.245/0.246 |
| Shikra (Vicuna-7B) [shikra] | softmax | 0.314/0.307 | 0.327/0.337 | 0.237/0.241 | 0.322/0.336 | 0.198/0.201 | 0.640/0.661 | 0.324/0.332 |
[…]
| InstructBLIP (Vicuna-7B) [iblip] | softmax | 0.359/0.437 | 0.683/0.689 | 0.200/0.283 | 0.253/0.367 | 0.263/0.304 | 0.629/0.663 | 0.337/0.382 |
[…]
| mPLUG-Owl (LLaMA-7B) [mplugowl] | softmax | 0.409/0.427 | 0.634/0.644 | 0.241/0.271 | 0.437/0.487 | 0.148/0.180 | 0.687/0.711 | 0.466/0.486 |
[…]
| LLaMA-Adapter-V2 [llamaadapterv2] | softmax | 0.354/0.363 | 0.464/0.506 | 0.275/0.329 | 0.298/0.360 | 0.251/0.257 | 0.604/0.666 | 0.412/0.425 |
[…]
| LLaVA-v1 (Vicuna-13B) [llava] | softmax | 0.462/0.457 | 0.442/0.462 | 0.264/0.280 | 0.404/0.417 | 0.285/0.297 | 0.626/0.684 | 0.349/0.372 |
III-C1 MLLM Performance
To measure the assessment ability, we evaluate the performance of 20 open-source MLLMs on 7 IQA datasets that are with at least 1,000 images and 15 human ratings per image [itu]. The experimental results are illustrated in Table VI. a) Primarily, we notice that the majority of MLLMs are notably better than NIQE on non-natural circumstances (CGI, AIGC, artificial distortions), showing their potential towards general-purpose evaluators on a broader range of low-level appearances. b) We also notice that without explicit alignment with human opinions during training, the most excellent MLLM, Emu2-Chat (which is based on the heaviest LLM, LLaMA-33B), can already outperform CLIP-ViT-Large-14 by a large margin (25%). These results have demonstrated that, though most MLLMs are still based on CLIP as visual encoders, their high capacity in the strong language decoder can do help them perform much better on visual quality assessment even without any explicit training.
III-C2 Superiority of softmax
In this section, we quantitatively evaluate the correlation with human perception on a simple argmax strategy between goodbad and our proposed softmax strategy. In Table XII, we select 5 MLLMs of different architectures and confirm that for all IQA datasets, the more measurable softmax strategy predicts better than the argmax strategy, which degenerates into only two scores, 0 and 1. Though the result is generally expected, the experiments validate that MLLMs have quantitative assessment ability hidden behind their word outputs, and prove the effectiveness of our softmax-based IQA strategy.
| Dataset Type | In-the-wild | Generated | Artificial | Average | ||||
[…]
| InfiMM (Zephyr-7B) [InfiMM] | ||||||||
[…]
| finebad | 0.331/0.368 | 0.500/0.527 | 0.190/0.251 | 0.305/0.366 | 0.309/0.324 | 0.555/0.651 | 0.411/0.430 | 0.372/0.417 |
| highlow | 0.412/0.382 | 0.539/0.492 | 0.216/0.194 | 0.586/0.524 | 0.173/0.171 | 0.674/0.698 | 0.429/0.429 | 0.433/0.413 |
| good+highpoor+low | 0.475/0.492 | 0.589/0.583 | 0.249/0.253 | 0.582/0.578 | 0.198/0.201 | 0.697/0.750 | 0.454/0.456 | 0.463/0.473 |
| good+finepoor+bad | 0.463/0.502 | 0.591/0.613 | 0.255/0.299 | 0.488/0.530 | 0.272/0.287 | 0.675/0.749 | 0.479/0.467 | 0.460/0.493 |
| good+high+finepoor+low+bad | 0.496/0.533 | 0.605/0.618 | 0.266/0.291 | 0.569/0.593 | 0.239/0.248 | 0.708/0.768 | 0.492/0.490 | 0.482/0.506 |
| Emu2-Chat (LLaMA-33B) [emu2] | ||||||||
[…]
| finebad | 0.663/0.540 | 0.711/0.702 | 0.359/0.362 | 0.601/0.631 | 0.285/0.334 | 0.770/0.599 | 0.846/0.830 | 0.605/0.571 |
| highlow | 0.685/0.644 | 0.721/0.703 | 0.333/0.334 | 0.633/0.647 | 0.255/0.237 | 0.779/0.793 | 0.830/0.795 | 0.605/0.593 |
| good+highpoor+low | 0.696/0.732 | 0.744/0.721 | 0.341/0.320 | 0.656/0.671 | 0.307/0.347 | 0.775/0.796 | 0.841/0.794 | 0.622/0.625 |
| good+finepoor+bad | 0.674/0.678 | 0.731/0.735 | 0.360/0.356 | 0.632/0.654 | 0.298/0.343 | 0.771/0.743 | 0.847/0.830 | 0.616/0.619 |
| good+high+finepoor+low+bad | 0.694/0.712 | 0.732/0.738 | 0.363/0.366 | 0.644/0.613 | 0.321/0.342 | 0.779/0.772 | 0.844/0.820 | 0.625/0.623 |
| InternLM-XComposer-VL (InternLM) [xcomposer] | ||||||||
[…]
| finebad | 0.546/0.597 | 0.720/0.736 | 0.341/0.389 | 0.626/0.671 | 0.213/0.227 | 0.681/0.708 | 0.494/0.479 | 0.517/0.544 |
| highlow | 0.543/0.590 | 0.704/0.720 | 0.331/0.372 | 0.612/0.656 | 0.223/0.251 | 0.716/0.755 | 0.490/0.500 | 0.517/0.549 |
| good+highpoor+low | 0.564/0.613 | 0.723/0.743 | 0.354/0.405 | 0.621/0.676 | 0.238/0.264 | 0.734/0.775 | 0.522/0.546 | 0.537/0.575 |
| good+finepoor+bad | 0.573/0.626 | 0.735/0.755 | 0.366/0.420 | 0.629/0.687 | 0.236/0.260 | 0.732/0.771 | 0.531/0.551 | 0.543/0.581 |
| good+high+finepoor+low+bad | 0.571/0.621 | 0.728/0.748 | 0.360/0.410 | 0.629/0.683 | 0.236/0.261 | 0.734/0.773 | 0.521/0.538 | 0.540/0.576 |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | ||||||||
[…]
| finebad | 0.453/0.469 | 0.457/0.482 | 0.258/0.288 | 0.303/0.333 | 0.294/0.302 | 0.558/0.617 | 0.389/0.420 | 0.388/0.416 |
| highlow | 0.474/0.476 | 0.370/0.386 | 0.261/0.262 | 0.432/0.429 | 0.266/0.269 | 0.669/0.716 | 0.304/0.331 | 0.397/0.410 |
| good+highpoor+low | 0.491/0.491 | 0.416/0.436 | 0.293/0.300 | 0.696/0.751 | 0.413/0.416 | 0.298/0.304 | 0.359/0.389 | 0.424/0.441 |
| good+finepoor+bad | 0.482/0.482 | 0.461/0.485 | 0.300/0.320 | 0.644/0.708 | 0.339/0.357 | 0.327/0.336 | 0.425/0.451 | 0.425/0.449 |
| good+high+finepoor+low+bad | 0.512/0.513 | 0.443/0.465 | 0.303/0.315 | 0.408/0.415 | 0.318/0.324 | 0.697/0.752 | 0.392/0.421 | 0.439/0.458 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | ||||||||
[…]
| finebad | 0.449/0.487 | 0.583/0.597 | 0.316/0.360 | 0.466/0.513 | 0.349/0.365 | 0.650/0.749 | 0.425/0.437 | 0.463/0.501 |
| highlow | 0.456/0.482 | 0.529/0.553 | 0.286/0.306 | 0.489/0.513 | 0.276/0.284 | 0.683/0.752 | 0.316/0.331 | 0.434/0.460 |
| good+highpoor+low | 0.462/0.484 | 0.548/0.573 | 0.303/0.327 | 0.480/0.509 | 0.283/0.294 | 0.687/0.763 | 0.350/0.363 | 0.445/0.473 |
| good+finepoor+bad | 0.463/0.483 | 0.579/0.596 | 0.321/0.356 | 0.467/0.505 | 0.326/0.339 | 0.670/0.762 | 0.420/0.426 | 0.464/0.495 |
| good+high+finepoor+low+bad | 0.474/0.498 | 0.565/0.588 | 0.314/0.345 | 0.488/0.521 | 0.311/0.322 | 0.692/0.771 | 0.382/0.392 | 0.461/0.491 |
| Qwen-VL (QwenLM) [Qwen-VL] | ||||||||
[…]
| finebad | 0.467/0.507 | 0.352/0.365 | 0.205/0.238 | 0.451/0.472 | 0.188/0.185 | 0.599/0.627 | 0.354/0.378 | 0.374/0.396 |
| highlow | 0.531/0.578 | 0.626/0.616 | 0.281/0.290 | 0.574/0.560 | 0.286/0.314 | 0.637/0.692 | 0.332/0.344 | 0.467/0.485 |
| good+highpoor+low | 0.539/0.600 | 0.684/0.673 | 0.299/0.324 | 0.565/0.568 | 0.306/0.330 | 0.660/0.721 | 0.414/0.422 | 0.495/0.520 |
| good+finepoor+bad | 0.495/0.558 | 0.596/0.581 | 0.264/0.307 | 0.521/0.548 | 0.270/0.270 | 0.640/0.691 | 0.435/0.449 | 0.460/0.486 |
| good+high+finepoor+low+bad | 0.541/0.600 | 0.632/0.617 | 0.286/0.316 | 0.570/0.577 | 0.301/0.318 | 0.664/0.719 | 0.416/0.429 | 0.487/0.511 |
| mPLUG-Owl (LLaMA-7B) [mplugowl] | ||||||||
[…]
| finebad | 0.357/0.398 | 0.622/0.636 | 0.260/0.290 | 0.422/0.475 | 0.178/0.224 | 0.606/0.646 | 0.536/0.534 | 0.426/0.457 |
| highlow | 0.353/0.369 | 0.610/0.624 | 0.176/0.187 | 0.436/0.464 | 0.110/0.124 | 0.662/0.663 | 0.361/0.378 | 0.387/0.401 |
| good+highpoor+low | 0.382/0.402 | 0.626/0.642 | 0.208/0.228 | 0.446/0.483 | 0.125/0.144 | 0.684/0.697 | 0.409/0.432 | 0.411/0.432 |
| good+finepoor+bad | 0.403/0.430 | 0.635/0.645 | 0.260/0.292 | 0.444/0.493 | 0.172/0.213 | 0.664/0.694 | 0.525/0.527 | 0.443/0.471 |
| good+high+finepoor+low+bad | 0.395/0.421 | 0.633/0.647 | 0.233/0.258 | 0.455/0.496 | 0.147/0.173 | 0.685/0.704 | 0.463/0.483 | 0.430/0.455 |
III-C3 Prompt Ensemble Effectiveness
As shown in Table XIII, the prompt ensemble strategy (as proposed in Eq. 2) on top-7 MLLMs (i.e. Emu2-Chat, InternLM-XComposer-VL, QWen-VL, InfiMM, LLaVA-v1.5 (13B), mPLUG-Owl, and LLaVA-v1.5 (7B)) can lead to up to 5% accuracy improvement (in average 1.7%). We believe it is a useful boost technique to improve the performance of MLLMs on the IQA task. Nevertheless, we also notice that different MLLMs perform best with different specific prompt combos. For example, the good+finepoor+bad performs best on InternLM-XComposer-VL, but comes with reduced accuracy on QWen-VL compared with only goodpoor. While goodpoor is proved overall best single word pair (except Emu2-Chat and LLaVA-v1.5 (13B)) for the evaluation and shows stable results across MLLMs, we decide to keep the current strategy (using the goodpoor combo) in Q-Bench+.
IV Conclusion
In this research, we introduce Q-Bench+, a benchmark designed to evaluate the advancements of MLLMs in low-level visual abilities. We evaluate the MLLMs from three aspects: perception of low-level visual attributes, description of low-level visual content, and assessment of image quality. Additionally, acknowledging the importance of discerning differences and similarities in image pairs, our benchmark encompasses both single images and image pairs in the perception and description tasks. To evaluate these abilities, we have compiled two multi-modal benchmark datasets focused on low-level vision, and introduced a unified softmax-based method for quantitative image quality assessment (IQA) in MLLMs. Our findings demonstrate that some advanced MLLMs exhibit commendable low-level visual abilities even without specialized low-level training. However, there’s still a significant journey ahead before MLLMs can become fully reliable assistants in general low-level visual tasks. We hope that the insights gained from Q-Bench+ will spur further development in MLLMs, particularly in improving their perception and understanding of low-level visual elements.
a
V Appendix
V-A Subjective Experiment
A total of twenty experts, each with professional skills and extensive experience in photography, are invited to participate in the subjective labeling experiment of Q-Bench+. The subjective experiment takes place in a laboratory environment with standard indoor lighting. A Dell-4K monitor, which supports a resolution of , is used for displaying the interfaces. The screenshots of interfaces can be referred to in Fig. 6. Each expert annotates up to 30 images a day to avoid fatigue, and every annotation is carefully reviewed by at least three other experts before acceptance. In this way, we ensure the accuracy and rigor of the Q-Bench+ labels to the greatest extent possible. This, in turn, makes the performance testing capability of Q-Bench+ more precise and meaningful.
V-B Distortions and Other Low-level Attributes Enumeration
V-B1 Distortions
Blurs [lens blur (out-of-focus), motion blur, zoom blur, gaussian blur, glass blur], Noises [gaussian noise, speckle noise, pepper noise], Artifacts [compression artifact, transmission error], Exposure Issues [under-exposure, over-exposure], Miscellaneous Artificial Distortions [pixelate, color-diffusion, jitter, etc]
V-B2 Other low-level attributes
Color [color style, color vividity], Lighting [bright, dim], Composition [Symmetrical, Rule-of-Thirds], Visual Styles [animation, realism, computer-generated, AI-generated], Photographic Methods [background bokeh (shallow DOF), high contrast, motion blur (on fast-moving objects), etc]
V-C Multi-choice Question for Kosmos-2
While Kosmos-2 performs generally well on the description and assessment tasks, we notice that it is hardly capable of answering a multi-choice question with the general prompt form applicable for other methods, as follows:
How is the clarity of the image?
(Question) [IMAGE_TOKEN] (Image)
Choose between one of the following options:
A. High (Correct) B. Medium(Wrong) C. Low(Wrong)
For most situations (86%) in our primary sample test with the prompts above, Kosmos-2 will directly append a new candidate (e.g., D. Excellent or D. Very Low) answer instead of choosing one option among them, denoted as prompt failure. This might be because the language model of Kosmos-2 has smaller capacity (1B) than other MLLMs that are based on LLaMA/MPT (7B/13B).
Considering that the prompt failure is actually not directly related with low-level perception, we try different prompt engineering techniques to reduce the prompt failure rate, and finalize with a simple modification which can limit the prompt failure to less than 10% in our sample set, as follows:
How is the clarity of the image?
(Question) [IMAGE_TOKEN] (Image)
Choose between one of the following options:
A. High (Correct) B. Medium(Wrong) C. Low(Wrong)
#Answer:
Nevertheless, we are still not able to eliminate the prompt failures for Kosmos-2. Henceforth, to systematically remove the negative effect of prompt failures on multi-choice questions for Kosmos-2, we conduct a choice-free special setting for it, i.e. close-set inference, via ranking the perplexity of different answers and choose the answer with minimum generative loss:
How is the clarity of the image?
[IMAGE_TOKEN] #Answer: High
loss:7.43 ✓ Choose this.
How is the clarity of the image?
[IMAGE_TOKEN] #Answer: Medium
loss:7.56 ✗
How is the clarity of the image?
[IMAGE_TOKEN] #Answer: Low
loss:7.92 ✗
As shown in Table IX, perplexity-based close-set inference can notably improve results of Kosmos-2. Considering that it is still the MLLM with fewest parameters among the ten models, its results are decent at its model size. More importantly, they validate that our observation of the prompt failure is reasonable.
| Sub-categories | Question Types | Quadrants of Low-level Concerns | Overall | # | |||||
[…]
| Distortion | Other | ||||||||
| random guess | 50.00% | 28.18% | 33.30% | 37.54% | 38.49% | 38.70% | 36.50% | 37.87% | - |
[…]
| ⋆⋆Kosmos-2 (close-set) | 61.48% | 37.13% | 40.76% | 40.04% | 50.88% | 45.30% | 58.15% | 47.26% | ✓ |
| Model (variant) | all | correct answer: Yes | correct answer: No | mean | de-biased # |
| random guess | 50.00% | 50.00% | 50.00% | 50.00% | - |
[…]
| LLaVA-v1 (Vicuna-13B) | 57.10% | 60.29% | 51.66% | 55.97% | 6 |
| MiniGPT-4 (Vicuna-13B) | 57.56% | 70.00% | 37.38% | 53.69% | 8 |
| LLaMA-Adapter-V2 | 67.12% | 68.80% | 64.76% | 66.78% | 2 |
| InstructBLIP (Flan-T5-XL) | 68.67% | 80.14% | 50.23% | 65.19% | 4 |
| InstructBLIP (Vicuna-7B) | 71.40% | 84.32% | 50.47% | 67.39% | 1 |
| Otter-v1 (MPT-7B) | 57.74% | 70.14% | 37.38% | 53.76% | 7 |
| IDEFICS-Instruct (LLaMA-7B) | 59.74% | 88.65% | 13.09% | 50.87% | 9 |
| mPLUG-Owl (LLaMA-7B) | 69.31% | 95.82% | 26.67% | 61.25% | 5 |
V-D “Yes or No?”: How Biased are MLLMs?
In this section, we take a deeper analysis on the Yes-or-No judgment ability of MLLMs, that whether these models can get similar accuracy on questions that should be answered with Yes, as those should be replied as No. Sadly, we notice that all MLLMs have higher prediction accuracy on Yes-questions than No-questions, while some MLLMs are more very severe biased (e.g., IDEFICS-Instruct). Considering that our LLVisionQA dataset contains more (62%) Yes-questions than No-questions (38%) and may introduce biases while comparing different MLLMs, we further compute a de-biased accuracy for all these methods, as the mean value of the accuracies on two types of questions, and present the respective de-biased rank for all participating MLLMs, as listed in Table X. We hope this study on the biases and the de-biased results can provide a fairer comparison among them, as well as bring insights on the future improvements of MLLMs for low-level perception.
V-E Settings for GPT Evaluation of Perception
Given GPT’s inherent variability, identical prompts can yield non-definitive responses. To address the impact of such situations on our evaluation, we’ve implemented a 5-round voting strategy. Under this approach, we pose the same prompt as defined in the following templates five times, taking the popular votes of GPT’s answers to determine the final outcome. Our human analysis on a sample set confirms that the 5-round voting strategy improves GPT evaluation accuracy from 93.2% to 98.4%, reducing errors to only 1/4 compared with the single-round evaluation.
Prompt Templates for GPT Evaluation:
#System: You are a helpful assistant that grades answers related to image quality and aesthetics. There are a lot of special terms or keywords related to image processing and photography. You will pay attention to the context of ’quality evaluation’ when grading.
#User: Assuming you are a grader, you will now be provided with a question [question] and a set of options [options] with option [options[0]] being the correct answer. Additionally, there will be an answer [answer] provided by a respondent. Please determine whether the respondentś answer is correct considering the context of the question. Even if the word choice is not completely the same, you can decide based on the given options and see whether the one in the answer is close enough to the given correct answer, The result is 1 if the answer is correct and else the result is 0. Please only provide the result in the following format: Result:
V-F Comparing natural images and AI-generated images
There might be doubt that pairwise comparison between natural images (NIs) and AI-generated images (AIGIs) is unnecessary since the criteria for assessing quality differ between these two types of images is different. However, we believe it is essential to continue conducting comparisons between NIs and AIGIs for several reasons:
Common Technical Quality Issues: AIGIs are generated based on prior knowledge learned from NIs and, as such, they may inherit common technical issues such as noise, blur, and incorrect exposure. Therefore, examining and comparing the technical quality of NIs and AIGIs is meaningful, as it can guide and improve the technical quality of AIGIs.
Aesthetic Quality Comparability: AIGIs and NIs also share comparability in aesthetic quality, such as in composition, color adjustment, and style. Investigating the aesthetic quality comparison between AIGIs and NIs can enhance the corresponding capabilities of MLLMs, thus improving the aesthetic quality of AI-generated content. Furthermore, considering the greater aesthetic freedom and diversity of AIGIs, comparing them with NIs can also provide more comprehensive aesthetic guidance for the photography of NIs.
V-G Settings for GPT Evaluation of Description
The detailed input prompt template for GPT to evaluate the three dimensions is listed as follows:
#System: You are a helpful assistant.
Completeness. #User: Evaluate whether the description [MLLM_DESC] completely includes the low-level visual information in the reference description [GOLDEN_DESC].
Please rate score 2 for completely or almost completely including reference information, 0 for not including at all, 1 for including part of the information or similar description.
Please only provide the result in the following format: Score: -
Preciseness. #User Evaluate whether the description [MLLM_DESC] precisely reflects the reference description [GOLDEN_DESC].
Please rate score 2 for totally no controversial low-level description, 1 for less controversial low-level description than matched descrpition, and 0 for more controversial low-level description than matched description. Please only provide the result in the following format: Score: -
Relevance. #User: Evaluate whether the description [MLLM_DESC] is relevant to the low-level visual information, which may include blur, noise, exposure, artifact, color, lighting, focus, composition, etc.
Please rate score 2 for completely relevant, 1 for partly relevant, and 0 for totally irrelevant.
Please only provide the result in the following format: Score:
In the prompt template, the [MLLM_DESC] denotes the output description from MLLMs, and [GOLDEN_DESC] denotes the golden description in the LLDescribe+ dataset.
| Dataset Type | Prompt | In-the-wild | Generated | Artificial | Average | ||||
[…]
| InfiMM (Zephyr-7B) [InfiMM] | Prompt1 | 0.471/0.541 | 0.633/0.610 | 0.257/0.276 | 0.576/0.576 | 0.263/0.261 | 0.731/0.775 | 0.491/0.461 | 0.489/0.500 |
| Prompt2 | 0.496/0.533 | 0.605/0.618 | 0.266/0.291 | 0.569/0.593 | 0.239/0.248 | 0.708/0.768 | 0.492/0.490 | 0.482/0.506 | |
[…]
| Prompt2 | 0.694/0.712 | 0.732/0.738 | 0.363/0.366 | 0.644/0.613 | 0.321/0.342 | 0.779/0.772 | 0.844/0.820 | 0.625/0.623 | |
[…]
| Prompt2 | 0.571/0.621 | 0.728/0.748 | 0.360/0.410 | 0.629/0.683 | 0.236/0.261 | 0.734/0.773 | 0.521/0.538 | 0.540/0.576 | |
[…]
| Prompt2 | 0.512/0.513 | 0.443/0.465 | 0.303/0.315 | 0.408/0.415 | 0.318/0.324 | 0.697/0.752 | 0.392/0.421 | 0.439/0.458 | |
[…]
| Prompt2 | 0.474/0.498 | 0.565/0.588 | 0.314/0.345 | 0.488/0.521 | 0.311/0.322 | 0.692/0.771 | 0.382/0.392 | 0.461/0.491 | |
[…]
| Prompt2 | 0.541/0.600 | 0.632/0.617 | 0.286/0.316 | 0.570/0.577 | 0.301/0.318 | 0.664/0.719 | 0.416/0.429 | 0.487/0.511 | |
[…]
| Prompt2 | 0.395/0.421 | 0.633/0.647 | 0.233/0.258 | 0.455/0.496 | 0.147/0.173 | 0.685/0.704 | 0.463/0.483 | 0.430/0.455 | |
| Dataset Type | In-the-wild | Generated | Artificial | |||||
[…]
| Shikra (Vicuna-7B) [shikra] | argmax | 0.178/0.201 | 0.277/0.281 | 0.152/0.169 | 0.248/0.267 | 0.071/0.065 | 0.513/0.562 | 0.245/0.246 |
| Shikra (Vicuna-7B) [shikra] | softmax | 0.314/0.307 | 0.327/0.337 | 0.237/0.241 | 0.322/0.336 | 0.198/0.201 | 0.640/0.661 | 0.324/0.332 |
[…]
| InstructBLIP (Vicuna-7B) [iblip] | softmax | 0.359/0.437 | 0.683/0.689 | 0.200/0.283 | 0.253/0.367 | 0.263/0.304 | 0.629/0.663 | 0.337/0.382 |
[…]
| mPLUG-Owl (LLaMA-7B) [mplugowl] | softmax | 0.409/0.427 | 0.634/0.644 | 0.241/0.271 | 0.437/0.487 | 0.148/0.180 | 0.687/0.711 | 0.466/0.486 |
[…]
| LLaMA-Adapter-V2 [llamaadapterv2] | softmax | 0.354/0.363 | 0.464/0.506 | 0.275/0.329 | 0.298/0.360 | 0.251/0.257 | 0.604/0.666 | 0.412/0.425 |
[…]
| LLaVA-v1 (Vicuna-13B) [llava] | softmax | 0.462/0.457 | 0.442/0.462 | 0.264/0.280 | 0.404/0.417 | 0.285/0.297 | 0.626/0.684 | 0.349/0.372 |
| Dataset Type | In-the-wild | Generated | Artificial | Average | ||||
[…]
| InfiMM (Zephyr-7B) [InfiMM] | ||||||||
[…]
| finebad | 0.331/0.368 | 0.500/0.527 | 0.190/0.251 | 0.305/0.366 | 0.309/0.324 | 0.555/0.651 | 0.411/0.430 | 0.372/0.417 |
| highlow | 0.412/0.382 | 0.539/0.492 | 0.216/0.194 | 0.586/0.524 | 0.173/0.171 | 0.674/0.698 | 0.429/0.429 | 0.433/0.413 |
| good+highpoor+low | 0.475/0.492 | 0.589/0.583 | 0.249/0.253 | 0.582/0.578 | 0.198/0.201 | 0.697/0.750 | 0.454/0.456 | 0.463/0.473 |
| good+finepoor+bad | 0.463/0.502 | 0.591/0.613 | 0.255/0.299 | 0.488/0.530 | 0.272/0.287 | 0.675/0.749 | 0.479/0.467 | 0.460/0.493 |
| good+high+finepoor+low+bad | 0.496/0.533 | 0.605/0.618 | 0.266/0.291 | 0.569/0.593 | 0.239/0.248 | 0.708/0.768 | 0.492/0.490 | 0.482/0.506 |
| Emu2-Chat (LLaMA-33B) [emu2] | ||||||||
[…]
| finebad | 0.663/0.540 | 0.711/0.702 | 0.359/0.362 | 0.601/0.631 | 0.285/0.334 | 0.770/0.599 | 0.846/0.830 | 0.605/0.571 |
| highlow | 0.685/0.644 | 0.721/0.703 | 0.333/0.334 | 0.633/0.647 | 0.255/0.237 | 0.779/0.793 | 0.830/0.795 | 0.605/0.593 |
| good+highpoor+low | 0.696/0.732 | 0.744/0.721 | 0.341/0.320 | 0.656/0.671 | 0.307/0.347 | 0.775/0.796 | 0.841/0.794 | 0.622/0.625 |
| good+finepoor+bad | 0.674/0.678 | 0.731/0.735 | 0.360/0.356 | 0.632/0.654 | 0.298/0.343 | 0.771/0.743 | 0.847/0.830 | 0.616/0.619 |
| good+high+finepoor+low+bad | 0.694/0.712 | 0.732/0.738 | 0.363/0.366 | 0.644/0.613 | 0.321/0.342 | 0.779/0.772 | 0.844/0.820 | 0.625/0.623 |
| InternLM-XComposer-VL (InternLM) [xcomposer] | ||||||||
[…]
| finebad | 0.546/0.597 | 0.720/0.736 | 0.341/0.389 | 0.626/0.671 | 0.213/0.227 | 0.681/0.708 | 0.494/0.479 | 0.517/0.544 |
| highlow | 0.543/0.590 | 0.704/0.720 | 0.331/0.372 | 0.612/0.656 | 0.223/0.251 | 0.716/0.755 | 0.490/0.500 | 0.517/0.549 |
| good+highpoor+low | 0.564/0.613 | 0.723/0.743 | 0.354/0.405 | 0.621/0.676 | 0.238/0.264 | 0.734/0.775 | 0.522/0.546 | 0.537/0.575 |
| good+finepoor+bad | 0.573/0.626 | 0.735/0.755 | 0.366/0.420 | 0.629/0.687 | 0.236/0.260 | 0.732/0.771 | 0.531/0.551 | 0.543/0.581 |
| good+high+finepoor+low+bad | 0.571/0.621 | 0.728/0.748 | 0.360/0.410 | 0.629/0.683 | 0.236/0.261 | 0.734/0.773 | 0.521/0.538 | 0.540/0.576 |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | ||||||||
[…]
| finebad | 0.453/0.469 | 0.457/0.482 | 0.258/0.288 | 0.303/0.333 | 0.294/0.302 | 0.558/0.617 | 0.389/0.420 | 0.388/0.416 |
| highlow | 0.474/0.476 | 0.370/0.386 | 0.261/0.262 | 0.432/0.429 | 0.266/0.269 | 0.669/0.716 | 0.304/0.331 | 0.397/0.410 |
| good+highpoor+low | 0.491/0.491 | 0.416/0.436 | 0.293/0.300 | 0.696/0.751 | 0.413/0.416 | 0.298/0.304 | 0.359/0.389 | 0.424/0.441 |
| good+finepoor+bad | 0.482/0.482 | 0.461/0.485 | 0.300/0.320 | 0.644/0.708 | 0.339/0.357 | 0.327/0.336 | 0.425/0.451 | 0.425/0.449 |
| good+high+finepoor+low+bad | 0.512/0.513 | 0.443/0.465 | 0.303/0.315 | 0.408/0.415 | 0.318/0.324 | 0.697/0.752 | 0.392/0.421 | 0.439/0.458 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | ||||||||
[…]
| finebad | 0.449/0.487 | 0.583/0.597 | 0.316/0.360 | 0.466/0.513 | 0.349/0.365 | 0.650/0.749 | 0.425/0.437 | 0.463/0.501 |
| highlow | 0.456/0.482 | 0.529/0.553 | 0.286/0.306 | 0.489/0.513 | 0.276/0.284 | 0.683/0.752 | 0.316/0.331 | 0.434/0.460 |
| good+highpoor+low | 0.462/0.484 | 0.548/0.573 | 0.303/0.327 | 0.480/0.509 | 0.283/0.294 | 0.687/0.763 | 0.350/0.363 | 0.445/0.473 |
| good+finepoor+bad | 0.463/0.483 | 0.579/0.596 | 0.321/0.356 | 0.467/0.505 | 0.326/0.339 | 0.670/0.762 | 0.420/0.426 | 0.464/0.495 |
| good+high+finepoor+low+bad | 0.474/0.498 | 0.565/0.588 | 0.314/0.345 | 0.488/0.521 | 0.311/0.322 | 0.692/0.771 | 0.382/0.392 | 0.461/0.491 |
| Qwen-VL (QwenLM) [Qwen-VL] | ||||||||
[…]
| finebad | 0.467/0.507 | 0.352/0.365 | 0.205/0.238 | 0.451/0.472 | 0.188/0.185 | 0.599/0.627 | 0.354/0.378 | 0.374/0.396 |
| highlow | 0.531/0.578 | 0.626/0.616 | 0.281/0.290 | 0.574/0.560 | 0.286/0.314 | 0.637/0.692 | 0.332/0.344 | 0.467/0.485 |
| good+highpoor+low | 0.539/0.600 | 0.684/0.673 | 0.299/0.324 | 0.565/0.568 | 0.306/0.330 | 0.660/0.721 | 0.414/0.422 | 0.495/0.520 |
| good+finepoor+bad | 0.495/0.558 | 0.596/0.581 | 0.264/0.307 | 0.521/0.548 | 0.270/0.270 | 0.640/0.691 | 0.435/0.449 | 0.460/0.486 |
| good+high+finepoor+low+bad | 0.541/0.600 | 0.632/0.617 | 0.286/0.316 | 0.570/0.577 | 0.301/0.318 | 0.664/0.719 | 0.416/0.429 | 0.487/0.511 |
| mPLUG-Owl (LLaMA-7B) [mplugowl] | ||||||||
[…]
| finebad | 0.357/0.398 | 0.622/0.636 | 0.260/0.290 | 0.422/0.475 | 0.178/0.224 | 0.606/0.646 | 0.536/0.534 | 0.426/0.457 |
| highlow | 0.353/0.369 | 0.610/0.624 | 0.176/0.187 | 0.436/0.464 | 0.110/0.124 | 0.662/0.663 | 0.361/0.378 | 0.387/0.401 |
| good+highpoor+low | 0.382/0.402 | 0.626/0.642 | 0.208/0.228 | 0.446/0.483 | 0.125/0.144 | 0.684/0.697 | 0.409/0.432 | 0.411/0.432 |
| good+finepoor+bad | 0.403/0.430 | 0.635/0.645 | 0.260/0.292 | 0.444/0.493 | 0.172/0.213 | 0.664/0.694 | 0.525/0.527 | 0.443/0.471 |
| good+high+finepoor+low+bad | 0.395/0.421 | 0.633/0.647 | 0.233/0.258 | 0.455/0.496 | 0.147/0.173 | 0.685/0.704 | 0.463/0.483 | 0.430/0.455 |
V-H Evaluation Details for Assessment Ability
V-H1 Example pseudo code for MLLMs on IQA
In Algo. 1, we provide an example of how to evaluate image quality with MLLMs. The algorithm is simple with only 9 lines, and could be easily integrated with any new MLLMs (based on causal LLMs), so as to allow these models to quantitatively predict the quality of images.
V-H2 Detailed performance for softmax & prompt ensemble strategy
In this section, we present a detailed analysis of the performance of the softmax and prompt ensemble strategies. This comprehensive evaluation aims to provide a precise comparison and demonstrate the effectiveness of the proposed strategies. The softmax and argmax comparison is illustrated in Table XII while the prompt ensemble performance is exhibited in Table XIII.
V-H3 Enriching prompt set
We concur that incorporating more complex synonyms in the prompt ensemble strategy can help mitigate word bias and enable the prompt-ensemble strategy to more accurately target positive or negative quality embeddings. At the same time, we remain open to exploring different choices for prompt ensembles. The selection of terms in our paper represents a preliminary attempt, thus our choices are intentionally simplistic, aimed only at demonstrating the effectiveness of the prompt-ensemble strategy. Further enriching this selection is a valuable direction in prompt engineering, potentially unlocking the full assessment potential of the model. Therefore, we follow your suggestions by enriching the prompts with more complex words: For the positive prompts, we employ the set {satisfactory, excellent, perfect}, and for the negative prompts, the set {terrible, unsatisfactory, subpar} is utilized. The outcomes are depicted in Table XI. From this table, we observe that using complex prompts yields performance results comparable to those achieved with simple prompts. This suggests that the contributions of complex and simple prompts within the prompt-ensemble strategy are not significantly different.
| Dimensions | Single | Pair | Assessment | ||
[…]
| Without Q-Instruct | |||||
[…]
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 60.07% | 3.21 | 52.25% | 3.22 | 0.424/0.445 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 61.40% | 3.47 | 52.05% | 3.44 | 0.444/0.474 |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 64.35% | 4.21 | 51.11% | 3.51 | 0.541/0.581 |
| With Q-Instruct | |||||
[…]
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 67.42% | 3.82 | 53.17% | 3.41 | 0.691/0.722 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 70.43% | 4.00 | 54.44% | 3.57 | 0.649/0.677 |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 70.37% | 4.25 | 53.21% | 3.66 | 0.686/0.711 |
[…]
| Best Close-Source Performance | 74.10% | - | 78.07% | - | -/- |
V-H4 IQA evaluation strategy for CLIP-ViT-Large-14
In Tab. VI, we compare the IQA performance of MLLMs with CLIP-ViT-Large-14, the visual backbone of the majority of MLLMs. Attempting to understand whether the new language part (LLM) can do better than the original language part of CLIP, we try to compare between CLIP and MLLMs in a relatively aligned setting. Firstly, noticing that most MLLMs will resize images into as their input sizes, we align this setting on CLIP, and ignore the strategies as proposed by [clipiqa]. Secondly, same as the strategy on MLLMs, we also apply softmax pooling between good and poor, as in the CLIP’s zero-shot classification format: a photo of good quality and a photo of poor quality. Besides the two alignments, the quality scores of CLIP-ViT-Large-14 are obtained as follows:
where represents the cosine similarity calculation function.
V-H5 Special IQA settings for Flan-T5-based InstructBLIP
For InstructBLIP [iblip] (Flan-T5-XL), different from the majority of LLaMA-based (or MPT-based Otter-v1) MLLMs, the two top-frequency tokens are high (89%) and low (8%) instead of the common goodpoor. Henceforth, based on our motivation to only modify the argmax into softmax and follow the default top-frequency output tokens of MLLMs, we replace the probabilities of goodpoor into those of highlow in Eq. 1 for T5, defined as follows:
As validated in our experiments (Table XIII, the highlow pair generally predicts better than goodpoor on majority of databases. The better performance on MLLM-specific top-frequency tokens by side validates the effectiveness of our methodology for MLLMs on IQA.
V-H6 Special validation protocol for CGIQA-6K
The CGIQA-6K [zhang2023subjective] dataset contains two separate sub-sets which consist of 3,000 game images and 3,000 movie images respectively, with different instructions for human annotators during its subjective experiments. Therefore, we validate the MLLMs’ assessment performance on the two sub-sets individually and average the results for the final exhibition. The results of NIQE and CLIP-ViT-Large-14 are also obtained under the same protocol for a fair comparison.
V-I MLLMs Finetuned with Low-level Dataset
Additionally, we have proceeded to conduct the corresponding experiments separately to compare the performance of MLLMs that are fine-tuned on low-level multi-modal dataset. Q-Instruct [q-instruct] is the first large-scale, low-level multi-modal dataset available, prompting our decision to fine-tune four popular MLLMs (LLaVA-v1.5 (Vicuna-v1.5-7B), LLaVA-v1.5 (Vicuna-v1.5-13B), mPLUG-Owl2 (LLaMA-7B), and InternLM-XComposer-VL (InternLM).) using this dataset. The experimental results are presented in Table XIV. From the table, we observe that fine-tuning with the low-level multi-modal dataset significantly enhances performance on single image tasks (all finetuned MLLMs are superior to the best open-source performance), but only slightly improves performance on image pair tasks. This suggests that while the low-level multi-modal dataset helps MLLMs acquire better low-level knowledge, it does not sufficiently enhance their comparison abilities, resulting in limited performance gains for image pair tasks.
V-J Statement on Data Contamination
The Q-bench+ contains three tasks, where the first two tasks, (A1) perception and (A2) description, are evaluated with our own datasets proposed with the paper. For these two tasks, the questions, answers, or low-level descriptions in the two datasets are not seen by any existing MLLMs. Half of LLVisionQA+ (i.e. the test subset) and full of LLDescribe+ labels are kept private, to avoid being added to the training sets of any MLLMs. We hope that this measure will allow Q-bench+ to have long-term significance as an indicator of low-level visual abilities.
For the third task, (A3) assessment, the situation is a bit more complicated. For open-source models as tested, almost all of them have provided their technical reports, and as far as we know, no image quality assessment (IQA) dataset has participated in the multi-modality training stages of them. While text knowledge about image quality assessment should have been injected to them (e.g. a blurry image is a low quality image) during their pure-language training stages, we think this should not be regarded as data contamination for IQA, because the images cannot be seen by a language model. Instead, they are important knowledge for MLLMs to better link particular visual attributes (blur) to human opinions (quality), which motivates us to explore MLLMs for these tasks.